OpenStudy (anonymous):

What is the value of the expression? 43 • 4^−5 A. 4^−15 B. 4^−2 C. 4^2 D. 4^8

Nnesha (nnesha):

when we multiply same bases we should `add` exponents \[\huge\rm x^m \times x^n=x^{m+n}\]

OpenStudy (anonymous):

alright so do i add?

Nnesha (nnesha):

yes right just the exponents

OpenStudy (anonymous):

from the equation, theres only one exponent, so what do I add that with?

Nnesha (nnesha):

there are 2 you didn't type it correctly \[\huge\rm 4^3 \times 4^{-5}\]

OpenStudy (anonymous):

Ooh, my bad. Ok, is it -2?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

so the answer is B?

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

haha thank you :)

Nnesha (nnesha):

np :=)